OpenStudy (anonymous):

FIND TWO POSITIVE NUMBERS WHOSE SUM IS 16 AND WHOSE PRODUCT IS 55

OpenStudy (anonymous):

guess and check for this one i guess 10 and 6, but \(10\times 6=60\) so that is not right maybe try \(11\) and \(5\)

OpenStudy (anonymous):

yeah 11 and 5. i know that but i have to do this in algebraic form. that’s the part I’m confused

OpenStudy (nikato):

For more algebraic approach, Let your two numbers be x and y So x+y=16. And xy=55 Subtract x on both sides in the first equation and get y=16-x Substitute this into second equation x(16-x)=55 Now solve for x

OpenStudy (anonymous):

omg thanks got it :)